# LargeAD: Large-Scale Cross-Sensor Data Pretraining for Autonomous Driving

### v1

@@ 6 Conclusion
In this paper, we introduced LargeAD, a scalable and generalizable framework designed for large-scale pretraining across diverse LiDAR datasets. Our approach leverages vision foundation models (VFMs) to generate semantically enriched superpixels, aligning 2D image features with LiDAR point clouds for improved representation learning. By integrating VFM-assisted contrastive learning, superpoint temporal consistency, and multi-source data pretraining, our framework achieves state-of-the-art performance on multiple 3D scene understanding tasks, including LiDAR-based semantic segmentation and 3D object detection. Extensive experiments conducted on eleven diverse datasets highlight the effectiveness of our framework in both in-domain and out-of-domain scenarios. Our framework not only excels in downstream generalization but also demonstrates superior robustness under out-of-distribution conditions. The ablation study further validates the importance of our design choices, showcasing the significant impact of incorporating multiple datasets during pretraining and the benefits of each individual component of our framework. The results underline the potential of LargeAD to advance real-world autonomous driving applications by providing a more versatile and resilient model capable of adapting to various sensor configurations and driving environments. In future work, we aim to extend our approach to incorporate additional sensor modalities, such as radar and thermal imaging, further broadening the scope of cross-modal pretraining for autonomous systems.
Acknowledgments. This work is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221- 0012), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).

### v2

@@ 6 Conclusion
In this paper, we introduced LargeAD, a scalable and generalizable framework designed for large-scale pretraining across diverse LiDAR datasets. Our approach leverages vision foundation models (VFMs) to generate semantically enriched superpixels, aligning 2D image features with LiDAR point clouds for improved representation learning. By integrating a VFM-assisted contrastive learning objective, superpoint temporal consistency, and multi-source data pretraining, our framework achieves state-of-the-art performance on multiple 3D scene understanding tasks, including LiDAR-based semantic segmentation and 3D object detection. Extensive experiments conducted on eleven diverse datasets highlight the effectiveness of our framework in both in-domain and out-of-domain scenarios. LargeAD excels not only in downstream generalization but also demonstrates strong robustness under out-of-distribution conditions. The ablation study further validates the importance of our design choices, showcasing the significant impact of incorporating multiple datasets during pretraining and the benefits of each individual component within our framework. Our results underline the potential of LargeAD to support real-world autonomous driving by enabling versatile and resilient models that adapt to diverse sensor configurations and driving environments. In future work, we aim to extend our approach to incorporate additional sensor modalities, such as radar and thermal imaging, further broadening the scope of cross-modal pretraining for autonomous systems.
Acknowledgments. This work is supported by Shanghai Artificial Intelligence Laboratory. This work is affiliated to the PJ-PRJ24DMX001 program.
This work is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221-0012, MOE-T2EP20223-0002). This research is also supported by cash and in-kind funding from NTU S-Lab and industry partner(s). Lingdong Kong is supported by the Apple Scholars in AI/ML Ph.D. Fellowship program.